OpenStudy (anonymous):

Simplify (x - 2)2 + (y + 5)2 = 36

OpenStudy (anonymous):

the twos on the outside of paraenthesis are exponents

OpenStudy (anonymous):

y = 15-x

OpenStudy (anonymous):

Thanks but can u show how u got that?

OpenStudy (ajprincess):

(a+b)^2=a^2+2ab+b^2

OpenStudy (ajprincess):

(a-b)^2=a^2-2ab+b^2

OpenStudy (anonymous):

Can u show your work not just the formula?

OpenStudy (anonymous):

first of all u will take 2 common so it will be : 2 (x-2)+(y+5)=36 then u will divide by 2 : (x-2)+(y+5)=18 then take the brackets out and combine x-2+y+5=18 ,x+y+3=18 ,move the 3 to da other side x+y=15 ,y=15-x Or x=15-y __________ U got it or need more explanation ?

OpenStudy (anonymous):

do you mean to put this in general form?

OpenStudy (anonymous):

perfect thanks!

OpenStudy (anonymous):

you can't do that... you said the 2's outside the parenthesis are exponents...